OpenStudy (anonymous):

general solution of (dP/dr)= P^2

OpenStudy (nincompoop):

use power rule I am thinking that is perimeter and radius?

OpenStudy (anonymous):

I think the answer is P=(Ce^r)^(1/2) but I just need to be sure

OpenStudy (accessdenied):

You can check your answer by taking the derivative and seeing if your equation substituting your solution gets you an undeniably true equation (like 0 = 0 is always true).

OpenStudy (anonymous):

separate the variables and integrate \[\int\limits \frac{ dP }{ P^2 }=\int\limits dr\] \[\int\limits P ^{-2}dP=r+c,\frac{ -1 }{P }=r+c\]

OpenStudy (nincompoop):

okay, this is interest rate

OpenStudy (anonymous):

right I did that and I got \[\ln (P^2)=r\] then I have to solve for P so for that I got \[P^2= e^{r+c}\]

OpenStudy (anonymous):

which is also \[P^2= (e ^{r})(e^c)\]

OpenStudy (anonymous):

then (because of my teacher) I have to substitute C for e^c that makes the whole thing if \[ C=e^c \] then \[P^2=Ce^r\]

OpenStudy (accessdenied):

Can we back up? (I can see where you are going to get to the solution you found). Where did you get: \( \ln \left( P^2 \right) = r \) from? From separating the variables and integrating?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

actually ln(P^2)=r+c

OpenStudy (anonymous):

my bad. I always forget to add in the c

OpenStudy (accessdenied):

Ohh, I can see what you did. Careful, this is not always true: \( \displaystyle \int \frac{1}{x^n} \ dx = \ln x^n + c \) In fact, it only works for n=1. \( \displaystyle \int \frac{1}{x} \ dx = \ln x^1 + c \) In any other case, we use power rule on a negative power: \( \displaystyle \int \frac{1}{x^n} \ dx = \int x^{-n} \ dx = \frac{x^{-n+1}}{-n+1} + c \) Notice that this latter integral /only/ fails on n=1, where the denominator is 0.

OpenStudy (anonymous):

Errrr... so its \[(P^-2)/2= e ^{r+c}\]

OpenStudy (anonymous):

no its not even that its ((P^-2)/2)=r+c

OpenStudy (accessdenied):

Allmosttt, The power was originally -2. We have to add 1 to the power and divide that off.

OpenStudy (anonymous):

Sorry please be patient with me. So it's (P^-1/-1)

OpenStudy (accessdenied):

It's fine. :) and yes, there you go! P^(-1) / -1 = r + c is good now.

OpenStudy (anonymous):

aww I was so excited I thought I had the right answer. Now I have to rework the whole thing. Hang on.

OpenStudy (anonymous):

So the general solution of (dP/dr)=P^2 is P= 1/(r+c)

OpenStudy (accessdenied):

you missed a very small detail from P^(-1) / -1 = r + c

OpenStudy (anonymous):

the negative

OpenStudy (anonymous):

damn.

OpenStudy (anonymous):

ok its -1/(r+c)

OpenStudy (accessdenied):

yes. easy to miss but important. Don't feel bad though, we always make mistakes, and it is good to learn from those happy little accidents. :p

OpenStudy (anonymous):

Okay thanks! You have been so helpful otherwise I would have missed 20 points on a take home math test!

OpenStudy (accessdenied):

if you are ever uncertain, too, we always have the ability to check just by substituting back into the original equation. finding P^2 and dP/dr and substituting back into the original equation will always create a true equation in the end if P is correct.

OpenStudy (anonymous):

Okay thanks again!

OpenStudy (accessdenied):

glad to help! :)